# MIRA: Medical Time Series Foundation Model for Real-World Health Data

### v6

@@ Abstract
A foundation model for medical time series, pretrained on ethically approved clinical datasets, can substantially reduce annotation burdens, minimize the need for task-specific tuning, and promote reliable transferability across healthcare institutions, data modalities, and clinical tasks, especially in data-scarce or privacy-sensitive environments. However, existing generalist time series foundation models struggle to handle medical time series data due to their inherent challenges, including irregular intervals, heterogeneous sampling rates, and frequent missing values. To address these challenges, we introduce MIRA, a unified foundation model specifically designed for medical time series forecasting. MIRA incorporates a Continuous-Time Rotary Positional Encoding that enables fine-grained modeling of variable time intervals, a frequency-specific mixture-of-experts layer that routes computation across latent frequency regimes to further promote temporal specialization, and a Continuous Dynamics Extrapolation Block based on Neural ODE that models the continuous trajectory of latent states, enabling accurate forecasting at arbitrary target timestamps. Pretrained on a large-scale and diverse medical corpus comprising over 454 billion time points collect from publicly available datasets, MIRA achieves reductions in forecasting errors by an average of 8% and 6% in out-of-distribution and in-distribution scenarios, respectively, when compared to other zero-shot and fine-tuned baselines. We also introduce a comprehensive benchmark spanning multiple downstream clinical tasks, establishing a foundation for future research in medical time series modeling. Our code is available at MIRA.
@@ 4.1 Experiment Setup
Downstream Clinical Datasets. We evaluate MIRA on a diverse suite of real-world clinical and public health datasets, spanning multi-modal physiological signal analysis, critical care monitoring, ambulatory bio-signal tracking, epidemiological surveillance, and healthcare resource utilization. To systematically study temporal robustness, we group datasets into two categories: (1) inherently irregular datasets, which exhibit irregular sampling and natural missing values due to clinical or observational workflows. These are CinC 2012 [silva2012predicting] and CinC 2019 [reyna2020early]. Furthermore, we use (2) originally regular datasets, i.e.,MIT-BIH [moody2001impact], Johns Hopkins COVID-19 Dataset [dong2022johns] ,CDC Influenza Hospitalizations Admissions (CDC-IHA)  [^1] https://www.cdc.gov/flu-forecasting/data-vis/current-week.html, Heart Rate [tan2020monash] and illness [DBLP:conf/iclr/WuHLZ0L23], for which we simulate irregularity by randomly masking 30% of time points. The full list of datasets and statistics is provided in Appendix G.
Baselines. We compare MIRA against 13 state-of-the-art forecasting models, which we categorize into three groups: (i) zero-shot foundation models, including Time-MoE [DBLP:journals/corr/abs-2409-16040]  [^2] We excluded Time-MoE Ultra and Sundial from baselines as they are not open-sourced., Moirai [DBLP:conf/icml/WooLKXSS24], Moirai-MoE [DBLP:journals/corr/abs-2410-10469], Moment [DBLP:conf/icml/GoswamiSCCLD24], TimeGPT [DBLP:journals/corr/abs-2310-03589], Timer [DBLP:conf/icml/LiuZLH0L24], Lag-Llama [rasul2023lag], TimesFM [DBLP:conf/icml/DasKSZ24], and Chronos [DBLP:journals/corr/abs-2403-07815], which require inputs to be interpolated for evaluation; and (ii) full-shot forecasting models, including ContiFormer [DBLP:conf/nips/ChenRWFSL23], T-PatchGNN [DBLP:conf/icml/ZhangYL0024], Neural-CDE [DBLP:conf/nips/RubanovaCD19], and ODE-RNN [DBLP:journals/corr/abs-1907-03907], which are specialized for irregularly sampled time series and require task-specific training. (iii) Continue pre-trained zero-shot foundation models on medical corpora, including Time-MoE [DBLP:journals/corr/abs-2409-16040], Moirai [DBLP:conf/icml/WooLKXSS24] and Chronos [DBLP:journals/corr/abs-2403-07815] which performance best on zero-shot evaluation. Implementation details for all baselines are provided in Appendix H.
@@ 4.2 Performance on Out-of-distribution Forecasting
Objective. We evaluated seven unseen benchmarks excluded from pre-training corpora. For comparison, we fine-tuned the full-shot forecasting models on the training split of each benchmark, while the zero-shot foundation models were evaluated directly without any task-specific training or fine-tuning. To validate the effectiveness of our model architecture, we additionally pre-trained existing foundation models on the same medical corpus. Notably, CINC 2012 and CINC 2019 were excluded from this evaluation for foundation models, as applying regular-grid interpolation at the finest resolution results in over 98% of time steps being interpolated, leading to poor performance for all baselines.
@@ 4.3 Performance on In-distribution Forecasting
Table 3: Zero-shot forecasting performance on out-of-distribution datasets that are regularly sampled but contain missing values. Reported values are averaged across all prediction lengths. Lower RMSE and MAE indicate better predictions. Red: the best; Blue: the second best.
Models Zero-shot Ours Zero-shot Baseline
@@ Appendix G Downstream Tasks and Datasets
CinC 2019[reyna2020early] Derived from the PhysioNet 2019 Challenge, this dataset includes time series recorded during in-hospital stays, including variables like vital signs, lab test results, and intervention indicators. The data are highly sparse and irregular, reflecting realistic clinical workflows. We focus on continuous value prediction using sliding window evaluation, emphasizing temporal robustness.

### v7

@@ Abstract
A foundation model for medical time series, pretrained on ethically approved clinical datasets, can substantially reduce annotation burdens, minimize the need for task-specific tuning, and promote reliable transferability across healthcare institutions, data modalities, and clinical tasks, especially in data-scarce or privacy-sensitive environments. However, existing generalist time series foundation models struggle to handle medical time series data due to their inherent challenges, including irregular intervals, heterogeneous sampling rates, and frequent missing values. To address these challenges, we introduce MIRA, a unified foundation model specifically designed for medical time series forecasting. MIRA incorporates a Continuous-Time Rotary Positional Encoding that enables fine-grained modeling of variable time intervals, a frequency-specific mixture-of-experts layer that routes computation across latent frequency regimes to further promote temporal specialization, and a Continuous Dynamics Extrapolation Block based on Neural ODE that models the continuous trajectory of latent states, enabling accurate forecasting at arbitrary target timestamps. Pretrained on a large-scale and diverse medical corpus comprising over 454 billion time points collect from publicly available datasets, MIRA achieves reductions in forecasting errors by an average of 8% and 6% in out-of-distribution and in-distribution scenarios, respectively, when compared to other zero-shot and fine-tuned baselines. We also introduce a comprehensive benchmark spanning multiple downstream clinical tasks, establishing a foundation for future research in medical time series modeling. Our code is available at Microsoft/MIRA.
@@ 4.1 Experiment Setup
Downstream Datasets. We evaluate MIRA on a diverse suite of real-world clinical and public health datasets, spanning multi-modal physiological signal analysis, critical care monitoring, ambulatory bio-signal tracking, epidemiological surveillance, and healthcare resource utilization. To systematically study temporal robustness, we group datasets into two categories: (1) inherently irregular datasets, which exhibit irregular sampling and natural missing values due to clinical or observational workflows. This are CinC 2012 [silva2012predicting]. Furthermore, we use (2) originally regular datasets, i.e.,MIT-BIH [moody2001impact], Johns Hopkins COVID-19 Dataset [dong2022johns] ,CDC Influenza Hospitalizations Admissions (CDC-IHA)  [^1] https://www.cdc.gov/flu-forecasting/data-vis/current-week.html, Heart Rate [tan2020monash] and illness [DBLP:conf/iclr/WuHLZ0L23], for which we simulate irregularity by randomly masking 30% of time points. The full list of datasets and statistics is provided in Appendix G.
Baselines. We compare MIRA against 13 state-of-the-art forecasting models, which we categorize into three groups: (i) zero-shot foundation models, including Time-MoE [DBLP:journals/corr/abs-2409-16040]  [^2] We excluded Time-MoE Ultra and Sundial from baselines as they are not open-sourced at the moment., Moirai [DBLP:conf/icml/WooLKXSS24], Moirai-MoE [DBLP:journals/corr/abs-2410-10469], Moment [DBLP:conf/icml/GoswamiSCCLD24], TimeGPT [DBLP:journals/corr/abs-2310-03589], Timer [DBLP:conf/icml/LiuZLH0L24], Lag-Llama [rasul2023lag], TimesFM [DBLP:conf/icml/DasKSZ24], and Chronos [DBLP:journals/corr/abs-2403-07815], which require inputs to be interpolated for evaluation; and (ii) full-shot forecasting models, including ContiFormer [DBLP:conf/nips/ChenRWFSL23], T-PatchGNN [DBLP:conf/icml/ZhangYL0024], Neural-CDE [DBLP:conf/nips/RubanovaCD19], and ODE-RNN [DBLP:journals/corr/abs-1907-03907], which are specialized for irregularly sampled time series and require task-specific training. (iii) Continue pre-trained zero-shot foundation models on medical corpora, including Time-MoE [DBLP:journals/corr/abs-2409-16040], Moirai [DBLP:conf/icml/WooLKXSS24] and Chronos [DBLP:journals/corr/abs-2403-07815] which performance best on zero-shot evaluation. Implementation details for all baselines are provided in Appendix H.
@@ 4.2 Performance on Out-of-distribution Forecasting
Objective. We evaluated seven unseen benchmarks excluded from pre-training corpora. For comparison, we fine-tuned the full-shot forecasting models on the training split of each benchmark, while the zero-shot foundation models were evaluated directly without any task-specific training or fine-tuning. To validate the effectiveness of our model architecture, we additionally pre-trained existing foundation models on the same medical corpus. Notably, CINC 2012 was excluded from this evaluation for foundation models, as applying regular-grid interpolation at the finest resolution results in over 98% of time steps being interpolated, leading to poor performance for all baselines.
@@ 4.3 Performance on In-distribution Forecasting
Table 3: Zero-shot forecasting performance on in-distribution datasets that are regularly sampled but contain missing values. Reported values are averaged across all prediction lengths. Lower RMSE and MAE indicate better predictions. Red: the best; Blue: the second best.
Models Ours Baseline